UNVEILING THE SWEET CONFORMATIONS OF KETOHEXOSES

\underline{C. BERMUDEZ}, I. PENA, C. CABEZAS, A. M. DALY, S. MATA, J. L. ALONSO, Grupo de Espectroscopía Molecular (GEM), Edificio Quifima, Laboratorios de Espectroscopía y Bioespectroscopía, Parque Científico, Universidad de Valladolid, 47011 Valladolid, Spain.

The conformational behavior of ketohexoses D-Fructose, L-Sorbose, D-Tagatose and D-Psicose has been revealed from their rotational spectra. A broadband microwave spectrometer (CP-FTMW)^{a} combined with a laser ablation (LA) source^{b} has been used to rapidly acquire the rotational spectra in the 6 to 12 GHz frequency range. All observed species are stabilized by complicated intramolecular hydrogen-bonding networks. Structural motifs related to the sweetness of ketohexoses are revealed.

aG. G. Brown, B. C. Dian, K. O. Douglass, S. M. Geyer, S. T. Shipman, B. H. Pate, Rev. Sci. Instrum. 2008, 79, 053103.

bS. Mata, I. Peña, C. Cabezas, J. C. López, J. L. Alonso, J. Mol. Spectrosc. 2012, 280, 91.